# A heralded single-photon source based on superpositions of squeezed states

### v1

@@ III Comparison with a heralded single-photon source from the two-mode squeezed state
Figure 8: Plot of $P(1,1;r;-)$, $\tilde{P}(1,1;r)$, and ${\cal N}_{-}(r)P(1,1;r;-)/4$ as functions of $r$ which are represented as the thick solid red, thin solid blue, and thick dashed red curves, respectively. The graph of $\tilde{P}(1,1;r)$ reaches the maximum value $1/4$ for $r=\mbox{arcsech}(1/\sqrt{2})=0.881$ and reaches the minimum value 00 for $r=0$.
In Fig. 8, we plot $P(1,1;r;-)$, $\tilde{P}(1,1;r)$, and ${\cal N}_{-}(r)P(1,1;r;-)/4$ as a function of $r$. When $r=\mbox{\rm arcsech}(1/\sqrt{2})=0.881$, $\tilde{P}(1,1;r)$ achieves its maximum value of $1/4$. In the limit of $r\to 0$, $P(1,1;r;-)$ and $\tilde{P}(1,1;r)$ approach to $1/2$ and 00, respectively. The fact that $\tilde{P}(1,1;r)$ becomes zero in the limit of $r\to 0$ is a major defect for generation of the heralded single photon from the pure two-mode squeezed state. Under a practical condition where the squeezing parameter $r$ is small, we cannot generate the single photon efficiently. Thus, on condition of detecting $|-\alpha\rangle_{2}$ and $|1\rangle_{a}$, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ as a source of the heralded single photons. However, the actual probability of the emission of the single photons for $|r;-\rangle$ converges to zero as $r\to 0$.
Now if we prepare a factory where we can produce many light beams of $|r;-\rangle$ by the cross-Kerr nonlinear medium and the homodyne measurements, we obtain heralded single photons efficiently for small $r$ because of $\lim_{r\to 0}P(1,1;r;-)=0.5$. Moreover, looking at Fig. 3(b), we note that the efficiency $E(r;-)$ is close to unity for small $r$. Hence, if we construct the heralded single-photon source with $|r;-\rangle$, the emission of the single photons is nearly deterministic. Contrastingly, if we use the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ for the heralded single-photon source, we cannot obtain the emission of the single photons efficiently because $\lim_{r\to 0}\tilde{P}(1,1;r)=0$. In Fig. 8, we note that $\tilde{P}(1,1;r)<P(1,1;r;-)$ for $0\leq r\leq 2$. Thus, $|r;-\rangle_{a}$ is preferable to $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$. Although we always obtain the heralded single photon when we detect the heralding single photon for the two-mode squeezed state and its efficiency is exactly equal to unity, looking at Fig. 3(b), we observe that $E(r;-)$ is nearly and sufficiently equal to unity, as well. Thus, regarding the efficiencies, the superposition of oppositely squeezed states $|r;-\rangle$ is as useful as the pure two-mode squeezed state.
@@
$\tilde{Q}(r,\eta)=\frac{2\eta\tanh^{2}r}{2-\eta[1-\cosh(2r)]}.$ (23)
@@
$\tilde{Q}_{1}(r,\eta)=\eta\frac{\tanh^{2}r}{\cosh^{2}r}.$ (24)
Then, we can compute the efficiency as
$\tilde{E}(r,\eta)=\frac{\tilde{Q}_{1}(r,\eta)}{\tilde{Q}(r,\eta)}=\eta+(1-\eta)\cosh^{-2}r.$ (25)
In Figs. 6(a) and (b), we plot $\tilde{Q}_{1}(r,\eta)$ and $\tilde{E}(r,\eta)$ as functions of $r$ for $\eta=0.9$ while in Figs. 7(a) and (b), we plot $\tilde{Q}_{1}(r,\eta)$ and $\tilde{E}(r,\eta)$ as functions of $r$ and $\eta$. Figure 6(a) shows that $\tilde{Q}_{1}(r,\eta)$ is nearly equal to zero for small $r$. This characteristic is in contrast to our proposed method because $Q_{1}(r,\eta;-)$ is larger than $0.4$ for small $r$. Comparing $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ in Fig. 6(b), we cannot find a large difference between them for small $r$. For this point, both the two methods are comparable. Examining Fig. 7(a), we note that $Q_{1}(r,\eta;-)$ is larger than $\tilde{Q}_{1}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. Moreover, $\lim_{r\to 0}Q_{1}(r,\eta;-)>0$ and $\lim_{r\to 0}\tilde{Q}_{1}(r,\eta)=0$ hold. Thus, for most realistic cases, the superposition of oppositely squeezed states is more beneficial to the heralded single-photon emission than the pure two-mode squeezed state. Looking at Fig. 7(b), we note that $E(r,\eta;-)<\tilde{E}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. However, for larger $r$ and smaller $\eta$ the difference is small.
@@ IV The second-order intensity correlation functions
We plot $g^{(2)}(0)$ for $|r;-\rangle$ and $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ in Fig. 9. We observe that $g^{(2)}(0)$ is always less for $|r;-\rangle$ in the shown region. In particular for $|r;-\rangle$, we notice that $g^{(2)}(0)$ approaches zero as $r\to 0$ while for the two-mode squeezed state it is always greater than zero for $r\geq 0$. Thus, for $\eta=0.9$ and $0\leq r\leq 0.504$, the single-photon emitter realized by $|r;-\rangle$ has a better quality than the two-mode squeezed state. We can understand the reason why this difference emerges as follows. The state $|r;-\rangle$ is a superposition of photon-number states, $|2\rangle$, $|6\rangle$, $|10\rangle$, …. Thus, dividing the state $|r;-\rangle$ on two ports of a 50-50 beam splitter, the probability that the imperfect photon detector detects the state $|2\rangle_{a}$ is supressed. Thus, $g^{(2)}(0)$ hardly suffers from an imperfection caused by the term $\eta(1-\eta)|2\rangle_{a}{}_{a}\langle 2|$ included in the observable $\hat{M}$ given by Eq. (17). By contrast, the pure two-mode squeezed state is a superposition of the states, $|0\rangle_{a}|0\rangle_{b}$, $|1\rangle_{a}|1\rangle_{b}$, $|2\rangle_{a}|2\rangle_{b}$, …, so that it is vulnerable to the imperfect term $\eta(1-\eta)|2\rangle_{a}{}_{a}\langle 2|$. In Fig. 9(b), we plot $g^{(2)}(0)$ as a function of $r$ and $\eta$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. We observe that $g^{(2)}(0)$ for our odd superposition of squeezed states is always smaller than that associated with the two-mode squeezed state for small $r$. This fact implies that the heralded single-photon emitter realized by $|r;-\rangle$ is more robust than that realized by $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ under the condition that we must use the imperfect single-photon detector.
@@ V Conclusion
In this paper, we proposed a new method to generate a heralded single-photon source created from a superposition of oppositely squeezed states. We outlined the method for producing this state by injecting squeezed and coherent light beams onto a cross-Kerr nonlinear medium and estimated the probability and efficiency of the emission of the heralded single photons. To evaluate the quality of our proposed method, we used the pure two-mode squeezed state for comparison. When the single-photon detector detects the heralding signals inaccurately, our method shows better performances for the second-order intensity correlation function $g^{(2)}(0)$ than the method of the pure two-mode squeezed states. This is because $|r\rangle-|-r\rangle$ is a superposition of photon-number states, $|2\rangle$, $|6\rangle$, $|10\rangle$, … and the probability amplitude of $|n\rangle_{a}|2\rangle_{b}$ for $n=0,1,2,...$ is suppressed when we inject it into the 50-50 beam splitter. That is an important advantage of scheme. However the generation of the odd superposition of squeezed states is more difficult leading to a natural trade-off.

### v2

@@ III Comparison with a heralded single-photon source from the two-mode squeezed state
Figure 8: Plot of $P(1,1;r;-)$, $\tilde{P}(1,1;r)$, and ${\cal N}_{-}(r)P(1,1;r;-)/4$ as functions of $r$ which are represented as the thick solid red, thin solid blue, and thick dashed red curves, respectively. The graph of $\tilde{P}(1,1;r)$ reaches the maximum value $1/4$ for $r=\mbox{arcsech}(1/\sqrt{2})=0.881$ and reaches the minimum value 00 for $r=0$. We put a small frame of plots of ${\cal N}_{-}(r)P(1,1;r;-)/4$ and $\tilde{P}(1,1;r)$ for $0\leq r\leq 0.006$ in the main frame. These plots show that ${\cal N}_{-}(r)P(1,1;r;-)/4$ and $\tilde{P}(1,1;r)$ are larger than $4\times 10^{-6}$ for $r\geq 0.004$. Thus, the heralded single-photon sources of our scheme and the pure two-mode squeezed state are more efficient than the SPDC even for small squeezing parameter $r$.
In Fig. 8, we plot $P(1,1;r;-)$, $\tilde{P}(1,1;r)$, and ${\cal N}_{-}(r)P(1,1;r;-)/4$ as a function of $r$. When $r=\mbox{\rm arcsech}(1/\sqrt{2})=0.881$, $\tilde{P}(1,1;r)$ achieves its maximum value of $1/4$. In the limit of $r\to 0$, $P(1,1;r;-)$ and $\tilde{P}(1,1;r)$ approach $1/2$ and 00, respectively. In the limit of $r\to 0$, both ${\cal N}_{-}(r)P(1,1;r;-)/4$ and $\tilde{P}(1,1;r)$ approach zero, sharing the same defect. However, in a practical circumstance where the squeezing parameter $r$ is small, our scheme with the postselection of measuring $|-\alpha\rangle_{2}$ shown in Fig. 1 gives the probability $P(1,1;r;-)$ to be nearly equal to $1/2$ around $r=0$. Hence, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ as a source of the heralded single photons.
Now if we prepare a factory where we can produce many light beams of $|r;-\rangle$ by the cross-Kerr nonlinear medium and the homodyne measurements, we obtain heralded single photons efficiently for small $r$ because of $\lim_{r\to 0}P(1,1;r;-)=0.5$. Moreover, looking at Fig. 3(b), we note that the conditional probability $P_{\mbox{\scriptsize c}}(r;-)$ is close to unity for small $r$. Hence, if we construct the heralded single-photon source with $|r;-\rangle$, the emission of the single photons is nearly deterministic. Contrastingly, if we use the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ for the heralded single-photon source, we cannot obtain the emission of the single photons efficiently because $\lim_{r\to 0}\tilde{P}(1,1;r)=0$. In Fig. 8, we note that $\tilde{P}(1,1;r)<P(1,1;r;-)$ for $0\leq r\leq 2$. Thus, $|r;-\rangle_{a}$ is preferable to $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$. Although we always obtain the heralded single photon when we detect the heralding single photon for the two-mode squeezed state and its conditional probability is exactly equal to unity, looking at Fig. 3(b), we observe that $P_{\mbox{\scriptsize c}}(r;-)$ is nearly and sufficiently equal to unity, as well. Thus, regarding the conditional probabilities, the superposition of oppositely squeezed states $|r;-\rangle$ is as useful as the pure two-mode squeezed state.
@@
$\tilde{P}_{\mbox{\scriptsize click}}(r,\eta)=\frac{2\eta\tanh^{2}r}{2-\eta[1-\cosh(2r)]}.$ (24)
@@
$\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)=\eta\frac{\tanh^{2}r}{\cosh^{2}r}.$ (25)
Then, we can compute the conditional probability as
$\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)=\frac{\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)}{\tilde{P}_{\mbox{\scriptsize click}}(r,\eta)}=\eta+(1-\eta)\cosh^{-2}r.$ (26)
In Figs. 6(a) and 6(b), we plot $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ as functions of $r$ for $\eta=0.9$ while in Figs. 7(a) and 7(b), we plot $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ as functions of $r$ and $\eta$. Figure 6(a) shows that $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ is nearly equal to zero for small $r$. This characteristic is in contrast to our proposed method because $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ is larger than $0.4$ for small $r$. Comparing $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ in Fig. 6(b), we cannot find a large difference between them for small $r$. For this point, both methods are comparable. Examining Fig. 7(a), we note that $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ is larger than $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. Moreover, $\lim_{r\to 0}P_{\mbox{\scriptsize click},1}(r,\eta;-)>0$ and $\lim_{r\to 0}\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)=0$ hold. Thus, for most realistic cases, the superposition of oppositely squeezed states is more beneficial to the heralded single-photon emission than the pure two-mode squeezed state. Looking at Fig. 7(b), we note that $P_{\mbox{\scriptsize click,c}}(r,\eta;-)<\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. However, for larger $r$ and smaller $\eta$ the difference is small.
@@ IV The second-order intensity correlation functions
We plot $g^{(2)}(0)$ for $|r;-\rangle$ and $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ in Fig. 9. We observe that $g^{(2)}(0)$ is always less for $|r;-\rangle$ in the shown region. In particular for $|r;-\rangle$, we notice that $g^{(2)}(0)$ approaches zero as $r\to 0$, while for the two-mode squeezed state it is always greater than zero for $r\geq 0$. Thus, for $\eta=0.9$ and $0\leq r\leq 0.504$, the single-photon emitter realized by $|r;-\rangle$ has a better quality than the two-mode squeezed state. We can understand the reason why this difference emerges as follows. The state $|r;-\rangle$ is a superposition of photon-number states, $|2\rangle$, $|6\rangle$, $|10\rangle$, …. Thus, dividing the state $|r;-\rangle$ on two ports of a 50-50 beam splitter, the probability that the imperfect photon detector detects the state $|2\rangle_{a}$ is suppressed. Thus, $g^{(2)}(0)$ hardly suffers from an imperfection caused by the term $\eta(1-\eta)|2\rangle_{a}{}_{a}\langle 2|$ included in the observable $\hat{M}$ given by Eq. (18). By contrast, the pure two-mode squeezed state is a superposition of the states, $|0\rangle_{a}|0\rangle_{b}$, $|1\rangle_{a}|1\rangle_{b}$, $|2\rangle_{a}|2\rangle_{b}$, …, so that it is vulnerable to the imperfect term $\eta(1-\eta)|2\rangle_{a}{}_{a}\langle 2|$. In Fig. 9(b), we plot $g^{(2)}(0)$ as a function of $r$ and $\eta$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. We observe that $g^{(2)}(0)$ for our odd superposition of squeezed states is always smaller than that associated with the two-mode squeezed state for small $r$. This fact implies that the heralded single-photon emitter realized by $|r;-\rangle$ is more robust than that realized by $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ under the condition that we must use the imperfect single-photon detector. In Fig. 9(a), we note that $g^{(2)}(0)$ of the pure two-mode squeezed state exhibits partial stability around $g^{(2)}(0)=0.22$ for $0\leq r\leq 1.0$. In contrast, our proposed method can achieve a high quality of single photons for the small $r$ regime ($0\leq r\leq 0.504$). This regime of the squeezing parameter $r$ is most relevant in the current experiment.
To compare our scheme with other single-photon sources, the method of the pure two-mode squeezed state in particular, we also need to count the probability of single-photon generation as shown in Fig. 3. When $r$ is smaller than $0.504$, the success probability of obtaining a single photon after the postselection is nearly equal to $1/2$, as shown by $P(1,1;r;-)$ in Fig. 3(b), hence both the success probability and the quality are high enough. However, as Fig. 3(a) indicates, the yielding of the success case is rather low when $r$ approaches zero. The usual trade-off between purity and overall efficiency remains in our scheme [44, 45].
@@ V Conclusion
In this paper, we proposed a method to generate a heralded single-photon source created from a superposition of oppositely squeezed states. We outlined the method for producing this state by injecting squeezed and coherent light beams onto a cross-Kerr nonlinear medium and estimated the probability and conditional probability of the emission of the heralded single photons. To evaluate the quality of our proposed method, we used the pure two-mode squeezed state for comparison. When the single-photon detector detects the heralding signals inaccurately, our method shows better performances for the second-order intensity correlation function $g^{(2)}(0)$ than the method of the pure two-mode squeezed states for small $r$ as explained in Sec. IV. This is because $|r\rangle-|-r\rangle$ is a superposition of photon-number states, $|2\rangle$, $|6\rangle$, $|10\rangle$, … and the probability amplitude of $|n\rangle_{a}|2\rangle_{b}$ for $n=0,1,2,...$ is suppressed when we inject it into the 50-50 beam splitter. That is an important advantage of our scheme. However the generation of the odd superposition of squeezed states is more difficult leading to a natural trade-off.